Solve the following homogeneous wave equation: un (2. t) = 4uxx(x, t), u(0.t) = u(t) = 0, u(x,0) = 0, (3,0) = 1.
Answer #1

20 DATE: 7 PAGE NO ut - Ullan Solution 4(0,67= uer, 67 = 0 U101,0)=0 By variable reparable method hey solution 42,1)= X(T) ThPAGE No Noco X(0730 = XA=0 G = 6-0 so No sowion Consider so that Gernt Gevine condi? X(0)=0 is satisfied iffa 20 to XOFC) -o20 DATE: PAGE No le cout) sinne la cod (207) thy sin (226) n=1 Note lenorzo as sinna n=1 n=1 44.(90) = + = an by sin (nu) But20 DATE: PAGE No So the final answer is. Unt- singura sim (2017) n=1 where by rinz (1-1-11 n = 12 -